Two marbles are drawn one after the other with replacement from a bag that contains:
9 red marbles, 8 blue marbles, and 3 green marbles.
Find the probability of the following events.
(a) One red marble and one green marble are drawn in either order.
The probability is :
(b) Two blue marbles are drawn.
The probability is :
(c) At least one marble is drawn that is not blue.
The probability is :
(d) The second marble is green if we know that the first marble is green.
The probability is:

Answer by Boreal(15235)

Two ways RG GR. 9/20*3/19 or 3/20*9/19. it is 54/380=0.1421
BB is 8/20*7/19=56/380=0.1474
For at least one marble, it is 1-0.1474=0.8526, since the only chice that doesn't work are both blues.
If the first marble is green, the second has 2/19 probability or 0.1053